Grade 9General Physics

A body weighs 64 kg on the surface of the earth. What is the gravitational force on it due to the earth at a height to one-third the radius of the earth??

Manju
10 Years agoGrade 9

1 Answer

Aqsa Ahmed
10 Years ago
SOLUTION :GIVEN- MASS=64 KG
GRAVITATIONAL FORCE=?
HEIGHT= ONE- THIRD THE RADIUS OF THE EARTH
AS ,
gh=g R2/(R+h)2
mg=mg* R2/(R+h)2
=64*R2/(R+R/3)2=64*R2/(16 R2/9)
=64*9/16
=4*9
=36kg